OpenStudy (anonymous):

There are three consecutive integers. The product of the smaller two integers added to 6 is equal to the product of the smallest and largest integer. What is the largest integer?

OpenStudy (anonymous):

so the two smaller integers add up to 6, and the smallest and largest add up to 6 also?

OpenStudy (anonymous):

a+b= 6 and a+c= 6,

OpenStudy (anonymous):

no the 2 smaller ones added to 6 equal the product of the smaller and larger numbers.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so a+b+6=a+c

OpenStudy (anonymous):

im sorry but i dont know how to do this one

OpenStudy (anonymous):

you good

OpenStudy (anonymous):

Let's solve your equation step-by-step. (x−1)(x)+6=(x−1)(x+1) Step 1: Simplify both sides of the equation. x2−x+6=x2−1 Step 2: Subtract x^2 from both sides. x2−x+6−x2=x2−1−x2 −x+6=−1 Step 3: Subtract 6 from both sides. −x+6−6=−1−6 −x=−7 Step 4: Divide both sides by -1. −x−1=−7−1 x=7 Answer: x=7

OpenStudy (anonymous):

the largest integer is 7